Question:
Solve the system of equations by elimination. Will a graph of the system have parallel, intersecting, or coinciding lines? How do you know?
4x + 3y = 5
2x − 5y = − 17

Answer by Alan3354(69443)

Solve the system of equations by elimination. Will a graph of the system have parallel, intersecting, or coinciding lines? How do you know?
4x + 3y = 5
2x - 5y = -17
-----
4x + 3y = 5
4x -10y = -34 --- 2nd eqn times 2
----------------- Subtract
13y = 39
y = 3
============
x = -1
----------------------------------
Will a graph of the system have parallel, intersecting, or coinciding lines?
Intersecting.
=========
How do you know?
If there's a unique solution, they intersect.
Intersection is (-1,3)
------------
If the 2 graphs have different slopes, they intersect, also.